American Health Systems currently has 6,400,000 shares of stock outstanding and will report earnings of $10 milion in the current year. The company is considering the issuance of 1,700,000 additional shares that will net $30 per share to the corporation. a. What ls the immediate dilution potential for this new stock issue? (Do not round intermediate calculations and round your answer to 2 decimal places.) ints Dilution per share b-1. Assume that American Health Systems can ean 9 percent on the proceeds of the stock issue in time to include them in the current years results. Calculate earnings per share. (Do not round intermediate calculations and round your answer to 2 decimal places.) Print Raferences Earnings per share b-2. Should the new issue be undertaken based on earnings per share? Yes O No Prex 1of 3l Next
Answer #1

a. Earnings per share before stock issue = $10,000,000/6,400,000 = $1.56

Earnings per share after stock issue = $10,000,000/8,100,000 = $1.23

Dilution = $1.56 - $1.23 = $0.33 per share

b.-1 Net income = $10,000,000 + .09(1,700,000 * $30)

Net income = $10,000,000 + .09($51,000,000)

Net income = $10,000,000 + $4,590,000

Net income = $14,590,000

Earnings per share after additional income EPS = $14,590,000 / 8,100,000 = $1.80

b.-2 Yes